OpenStudy (anonymous):

PLEASE HELP I'VE BEEN DOING THIS FOR 2 HOURS!! The table shows a meteorologist's predicted temperatures for an April day in Washington D.C. Use quadratic regression to find a quadratic model for this data. (Use the 24-hour clock to represent times after noon.) Time Predicted Temperature 8 a.m. 50.68 10 a.m. 62.24 12 p.m.. 69.75 2 p.m.. 73.2 4 p.m... 72.59 6 p.m.. 67.93 a.) –0.007x2 + 15.907x – 40.131 b.) –0.507x2 + 14.907x – 36.131 c.) –0.507x2 – 14.907x + 36.131 d.) –1.007x2 + 13.907x – 32.131

OpenStudy (anonymous):

@satellite73 helppp!

OpenStudy (anonymous):

@Algebraic! can you help me?? pleeeease?

OpenStudy (anonymous):

@Zarkon helpppp!!

OpenStudy (anonymous):

Refer to the attachment. Calculations provided by Mathematica 8 Home Edition.

OpenStudy (anonymous):

1. y=3x^2-2x+2 2. y=4x^2+5x-6 3. P(x)=33x^2-24x+635; 2555 waterfowl 4. -0.507x^2+14.907x-36.131 5. 0 < or = to x < or = to 8